OpenStudy (anonymous):

10. Write the equation of a hyperbola with vertices (0, -2) and (0, 2) and co-vertices (-4, 0) and (4, 0). MEDAL WILL BE GIVEN ! (づ ̄ ³ ̄)づ

OpenStudy (anonymous):

i think you have it, give me one second to check

OpenStudy (anonymous):

oh hold on maybe not

OpenStudy (anonymous):

oh ok lets draw a picture

OpenStudy (anonymous):



OpenStudy (anonymous):

wow damned ugly picture, but that is from the info given verices at \((0,2)\) and \((0,-2)\) from the orientation, can you tell what comes first, the (x^2\) or the \(y^2\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and what number goes underneath the \(y^2\) ?

OpenStudy (anonymous):

it is 4 because it crosses at 2 and \(2^2=4\)

OpenStudy (anonymous):

so you know it is \[\frac{y^2}{4}-\frac{x^2}{b^2}=1\] now how about \(b\) ?

OpenStudy (anonymous):

hold on, don't forget you know the co vertices are \((-4,0)\) and \((0,4)\) use that to find the number that goes under the \(x^2\) term

OpenStudy (anonymous):

yeah, well i got froze out for a while anyway, can you see what goes under the \(x^2\) term? don't forget it is a square

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{y^2}{4}-\frac{x^2}{16}=1\]